# Hallucination of Multimodal Large Language Models: A Survey

### v1

@@ 5.1. Data
As discussed in the section on hallucination causes 3, data is one of the primary factors inducing hallucination in MLLMs. For mitigating hallucination, recent works make attempts on data, including introducing negative data (Liu et al., 2023d), introducing counterfactual data (Yu et al., 2023b), and reducing noise and errors in existing dataset (Wang et al., 2023a; Yue et al., 2024).
@@
ReCaption (Wang et al., 2023a) This work proposes a framework called ReCaption to rewrite the text captions of existing image-text pairs in datasets. The framework comprises two steps: 1) keyword extraction, which extracts verbs, nouns, and adjectives from the caption; and 2) caption generation, which employs an LLM to generate sentences based on the extracted keywords. Ultimately, the framework produces a set of high-quality image-caption pairs. Experiment results show that the model trained on the rewritten caption dataset has higher accuracy on certain benchmarks, such as the POPE benchmark (Li et al., 2023a). Despite the performance improvement, the question of why rewritten captions can reduce hallucination remains an open problem.
@@ 5.2.1. Scale-up Resolution
Enhancing the perception ability of MLLMs has been shown to improve their overall performance and reduce hallucination (Liu et al., 2023c, b; Zhai et al., 2023; Chen et al., 2023a). One important update when upgrading from LLaVA (Liu et al., 2023c) to LLaVA-1.5 (Liu et al., 2023b) is to scale up the CLIP ViT vision encoder from CLIP-ViT-L-224 to CLIP-ViT-L-336, resulting in considerable performance improvement. Qwen-VL (Bai et al., 2023) has shown the effectiveness of gradually enlarging image resolution from $224\times 224$ to $448\times 448$. InternVL (Bai et al., 2023) scales up the vision encoder to 6 billion parameters, enabling processing of high-resolution images. Regarding hallucination, HallE-Switch (Zhai et al., 2023) has investigated the impact of vision encoder resolution on its proposed CCEval benchmark. Among the three studied vision encoders (CLIP-ViT-L-112, CLIP-ViT-L-224, CLIP-ViT-L-336), higher resolution generally results in lower degrees of hallucination. These works indicate that scaling up vision resolution is a straightforward yet effective solution.
@@ 5.2.2. Versatile Vision Encoders
Several studies (Tong et al., 2024; He et al., 2024; Jain et al., 2023) have investigated vision encoders for MLLMs. Typically, the CLIP ViT image encoder is used as the vision encoder in most MLLMs thanks to its remarkable ability to extract semantic-rich features. However, CLIP has been shown to lose some visual details compared to pure vision models like DINO ViT (Caron et al., 2021). Therefore, recent studies have proposed complementing this information loss by incorporating visual features from other vision encoders. The work of (Tong et al., 2024) proposes mixing features from CLIP ViT and DINO ViT. Specifically, it experimented with additive and interleaved features. Both settings show that there is a trade-off between the two types of features. A more dedicated mechanism is needed.
Concurrently, a visual expert-based model proposed in (He et al., 2024) aims to mitigate the information loss caused by the CLIP image encoder. Instead of merely mixing features, this paper enhances the visual perception ability of MLLMs by focusing on knowledge enhancement, relying on two pivotal modules: multi-task encoders and the structural knowledge enhancement module. The multi-task encoders are dedicated to integrating various types of latent visual information extracted by multiple visual encoders. Additionally, the structural knowledge enhancement module is designed to utilize visual tools, such as OCR tools and object detectors, to extract prior knowledge from visual inputs.
Following the approach of the structural knowledge enhancement module in (He et al., 2024), another line of research investigates the utilization of vision tool models to enhance the perception of MLLMs. VCoder (Jain et al., 2023) utilizes additional perception formats, such as segmentation masks and depth maps, to enhance the object identification ability of the MLLM. Another work (Jiao et al., 2024) ensembles additional object detection and optical-character recognition models into the MLLM architecture. It also explores various ways to integrate this information, including training-free infusion, LoRA (Hu et al., 2021) augmented retraining, and LoRA augmented finetuning.
@@ 5.3.1. Auxiliary supervision
Accordingly, the work of (Chen et al., 2023c) constructs a fine-grained vision instruction dataset based on Panoptic Scene Graph (PSG), called Relation-Associated Instruction (RAI-30k). In addition to standard dialogues, each instruction in RAI-30k is associated with a relation annotation in PSG, which includes mask annotations for related instances. With these additional annotations, it further supervises MLLMs with mask prediction loss using a state-of-the-art expert vision model, SAM (Kirillov et al., 2023), guiding MLLMs to focus on highly-related image content. With the additional supervision from the mask prediction loss, MLLMs are encouraged to extract features that can better represent these crucial instances, thus generating more accurate responses and mitigating vision hallucination. The intuitive idea of supervising MLLMs with grounding shows promising performance in mitigating hallucination.
Another line of work analyzes the training loss from the perspective of embedding space distribution. As introduced earlier, popular MLLMs typically project the encoded vision features into the input space of a specific LLM. A recent work, HACL (Jiang et al., 2023), argues that an ideal projection should blend the distribution of visual and textual embeddings. However, despite visual projection, a significant modality gap exists between textual and visual tokens, suggesting that the current learned interfaces are not effective in mapping visual representations into the textual representation space of LLMs. This issue potentially exacerbates the tendency for MLLMs to generate more hallucinations. Therefore, HACL proposes enhancing the alignment between visual and textual representations through contrastive loss. Texts with hallucinations are used as hard negative examples for image anchors. The loss pulls representations of non-hallucinating text and visual samples closer while pushing representations of non-hallucinating and hallucinative text apart. Experiment results show that this method not only reduces hallucination but also enhances performance on other popular benchmarks.
@@ 5.3.2. Reinforcement Learning
A more recent work, POVID (Zhou et al., 2024), challenges the assumption underlying previous DPO-based methods. These methods rely on the traditional preference data generation process in LLMs, where both preferred and dispreferred responses may potentially be incorrect. Therefore, this work proposes the Preference Optimization in VLLM with AI-Generated Dispreferences (POVID) framework, aiming to exclusively generate dispreferred feedback data using AI models. The dispreferred data is generated by: 1) utilizing GPT-4V to introduce plausible hallucinations into the answer, and 2) provoking inherent hallucination by introducing noise into MLLMs. In the DPO optimization framework, the ground-truth multimodal instructions serves as the preferred answers.
@@
HalDetect (Gunjal et al., 2023) first introduces the M-HalDetect dataset for detecting hallucinations, which covers a wide range of hallucinatory content, including non-existent objects, unfaithful descriptions, and inaccurate relationships. It then proposes a multimodal reward model to detect hallucinations generated by MLLMs. The reward model is trained on the M-HalDetect dataset to identify hallucinations in the generated text. To utilize the trained reward model to reduce hallucinations, the authors introduced Fine-grained Direct Preference Optimization (FDPO). FDPO uses fine-grained preferences from individual examples to directly reduce hallucinations in generated text by enhancing the model’s ability to distinguish between accurate and inaccurate descriptions.
@@
Another similar work, ViGoR (Yan et al., 2024), also designs a fine-grained reward model to update pre-trained MLLMs, aiming to improve visual grounding and reduce hallucination. The reward modeling in this work encompasses both human preferences and automatic metrics. Specifically, it collects human judgment and preferences for the responses generated by MLLMs by asking crowd-workers to provide fine-grained feedback at the sentence level. The collected human preference data is used to train a reward model. Additionally, it leverages advanced vision perception models to automatically score the grounding and fidelity of the text generated by an MLLM. Both sources are combined into a single reward score during the reinforcement learning procedure.
@@ 5.3.3. Unlearning
Unlearning refers to a technique designed to induce a model to ’forget’ specific behaviors or data, primarily through the application of gradient ascent methods (Cao and Yang, 2015). Recently, unlearning for LLMs has been receiving increasing attention (Jang et al., 2022), effectively eliminating privacy vulnerabilities in LLMs. In the context of MLLMs, a recent work (Xing et al., 2024) introduces the Efficient Fine-grained Unlearning Framework (EFUF), applying an unlearning framework to address the hallucination problem. Specifically, it utilizes the CLIP model to construct a dataset comprised of both positive samples and negative (hallucinated) samples. The training loss is applied separately for positive and negative at the sub-sentence level. To the best of our knowledge, EFUF (Xing et al., 2024) is the first and only work that applies the unlearning framework to the task of hallucination mitigation, opening up a new path for future research.
@@ 5.4.1. Generation Intervention
VCD (Visual Contrastive Decoding) (Leng et al., 2023) is designed to suppress the statistical biases and language priors in MLLMs during the decoding phase. The main assumption of VCD is that a distorted visual input would lead to text responses with more biases and priors. Thus, by contrasting output distributions derived from original and distorted visual inputs, VCD aims to effectively reduce the over-reliance on statistical bias and language priors. Specifically, the decoding probability distribution is calibrated using the reference (distorted) distribution.
Following the same idea of contrastive decoding, IBD (Zhu et al., 2024) proposes an image-biased decoding strategy. Specifically, IBD involves computing a more reliable next-token probability distribution by contrasting the predictions of the original model with those of an image-biased model, which focuses more on the image information. The image-based model is created by modifying the attention weight matrix structure within the original model, without altering its parameters. This approach emphasizes the knowledge of the image-biased model and diminishes that of the original model, which may be text-biased. Thus, it encourages the extraction of correct content while suppressing hallucinations resulting from textual over-reliance.
@@
MARINE (Zhao et al., 2024) proposes a training-free approach. It employs an additional vision encoder for object grounding and utilizes the grounded objects to guide the decoding process. Specifically, it innovatively adapts the classifier-free guidance (Ho and Salimans, 2022) technique to implement guided decoding, showing promising performance in emphasizing the detected objects while reducing hallucination in the text response.
@@
HALC (Chen et al., 2024b) provides a key insight that when decoding a specific token in the MLLM, identifying a token-wise optimal visual context to provide the most informative visual grounding can effectively reduce hallucination. Visual context refers to the visual tokens that can be grounded from the generated text response. An oracle study showed that decoding from the provided optimal visual contexts eliminates over 84.5% of hallucinations. Based on the insight and observation, the authors designed mechanisms to locate the fine-grained visual information to correct each generated token that might be hallucinating. This is essentially a visual content-guided decoding strategy. In addition to token-level correction, HALC also incorporates a matching-based beam search that utilizes a visual matching score to steer the generation of the final outputs, balancing both object hallucination mitigation and text generation quality.
@@
The work of OPEAR (Huang et al., 2023b) makes an interesting observation that most hallucinations are closely tied to the knowledge aggregation patterns manifested in the self-attention matrix, i.e., MLLMs tend to generate new tokens by focusing on a few summary tokens rather than all the previous tokens. Such a partial over-trust inclination results in neglecting image tokens and describing the image content with hallucination. Based on this observation, a decoding method for MLLMs grounded in an Over-trust Penalty and a Retrospection-Allocation strategy is proposed. First, a penalty term on the model logits is introduced during the MLLM beam-search decoding process to mitigate the over-trust issue. Additionally, to handle the hard cases that cannot be addressed by the penalty term, a more aggressive strategy called the rollback strategy is proposed to retrospect the presence of summary tokens in the previously generated tokens and reallocate the token selection if necessary.

### v2

@@ 5.1. Data
As discussed in the section on hallucination causes 3, data is one of the primary factors inducing hallucination in MLLMs. For mitigating hallucination, recent works make attempts on data, including introducing negative data (Liu et al., 2023d), counterfactual data (Yu et al., 2023b; Chen et al., 2025a), reasoning data (Zhang et al., 2024c), and reducing noise and errors in existing dataset (Wang et al., 2024d; Yue et al., 2024).
5.1.1. Negative Data
@@ 5.1.2. Counterfactual Data
Recently, PerturboLLaVA (Chen et al., 2025a) explores a similar strategy. It reduces the model’s heavy dependence on the language prior by incorporating adversarially perturbed text during training. Specifically, it introduce carefully designed perturbations that aligns with the general knowledge but conflict with the visual content, intentionally misleading the model based on its language prior. This perturbative training enforces the model to scrutinize the image content when predicting every token, rather than hallucinating contents from the text hints. Essentially, this method adjusts the model’s conditional distribution to depend more heavily on the image and less on the perturbation text, which leads to more robust multimodal capability.
5.1.3. Reasoning Data
REVERIE (Zhang et al., 2024c) This paper argues that without intermediate reasoning steps, models may establish superficial shortcuts between instructions and responses, failing to internalize the inherent reasoning logic. To address this challenge, it propose reflective instruction tuning, which integrates rationale learning into visual instruction tuning. Unlike previous methods that learning from responses only, this approach entails the model predicting rationales justifying why responses are correct or incorrect. This instruction tuning is supported by a specially curated dataset REVERIE. Each instruction is meticulously annotated with a corresponding pair of correct and confusing responses, alongside comprehensive rationales elucidating the justification behind the correctness or erroneousness of each response.
5.1.4. Clean Data
ReCaption (Wang et al., 2024d) This work proposes a framework called ReCaption to rewrite the text captions of existing image-text pairs in datasets. The framework comprises two steps: 1) keyword extraction, which extracts verbs, nouns, and adjectives from the caption; and 2) caption generation, which employs an LLM to generate sentences based on the extracted keywords. Ultimately, the framework produces a set of high-quality image-caption pairs. Experiment results show that the model trained on the rewritten caption dataset has higher accuracy on certain benchmarks, such as the POPE benchmark (Li et al., 2023a). Despite the performance improvement, the question of why rewritten captions can reduce hallucination remains an open problem.
@@ 5.2.1. Scale-up Resolution
Enhancing the perception ability of MLLMs has been shown to improve their overall performance and reduce hallucination (Liu et al., 2023c, b; Zhai et al., 2023; Chen et al., 2024d). One important update when upgrading from LLaVA (Liu et al., 2023c) to LLaVA-1.5 (Liu et al., 2023b) is to scale up the CLIP ViT vision encoder from CLIP-ViT-L-224 to CLIP-ViT-L-336, resulting in considerable performance improvement. Qwen-VL (Bai et al., 2023) has shown the effectiveness of gradually enlarging image resolution from $224\times 224$ to $448\times 448$. InternVL (Chen et al., 2024d) scales up the vision encoder to 6 billion parameters, enabling processing of high-resolution images. Regarding hallucination, HallE-Switch (Zhai et al., 2023) has investigated the impact of vision encoder resolution on its proposed CCEval benchmark. Among the three studied vision encoders (CLIP-ViT-L-112, CLIP-ViT-L-224, CLIP-ViT-L-336), higher resolution generally results in lower degrees of hallucination. These works indicate that scaling up vision resolution is a straightforward yet effective solution.
@@ 5.2.2. Versatile Vision Encoders
Several studies (Tong et al., 2024; He et al., 2024a; Jain et al., 2023) have investigated vision encoders for MLLMs. Typically, the CLIP ViT image encoder is used as the vision encoder in most MLLMs thanks to its remarkable ability to extract semantic-rich features. However, CLIP has been shown to lose some visual details compared to pure vision models like DINO ViT (Caron et al., 2021). Therefore, recent studies have proposed complementing this information loss by incorporating visual features from other vision encoders. The work of (Tong et al., 2024) proposes mixing features from CLIP ViT and DINO ViT. Specifically, it experimented with additive and interleaved features. Both settings show that there is a trade-off between the two types of features. A more dedicated mechanism is needed.
Concurrently, a visual expert-based model proposed in (He et al., 2024a) aims to mitigate the information loss caused by the CLIP image encoder. Instead of merely mixing features, this paper enhances the visual perception ability of MLLMs by focusing on knowledge enhancement, relying on two pivotal modules: multi-task encoders and the structural knowledge enhancement module. The multi-task encoders are dedicated to integrating various types of latent visual information extracted by multiple visual encoders. Additionally, the structural knowledge enhancement module is designed to utilize visual tools, such as OCR tools and object detectors, to extract prior knowledge from visual inputs.
Following the approach of the structural knowledge enhancement module in (He et al., 2024a), another line of research investigates the utilization of vision tool models to enhance the perception of MLLMs. VCoder (Jain et al., 2023) utilizes additional perception formats, such as segmentation masks and depth maps, to enhance the object identification ability of the MLLM. Another work (Jiao et al., 2024) ensembles additional object detection and optical-character recognition models into the MLLM architecture. It also explores various ways to integrate this information, including training-free infusion, LoRA (Hu et al., 2021) augmented retraining, and LoRA augmented finetuning. EAGLE (Villa et al., 2025) retrains the vision encoder with enhanced visual grounding capability, reducing hallucination in downstream application on MLLMs.
@@ 5.2.3. Dedicated Module
In the work of PATCH (Shang et al., 2024), it analyze the hallucination problem from an architectural perspective. Correspondingly, it proposes a tuning strategy aligning visual and textual features in the semantic space to mitigate object hallucinations. Concretely, it inserts several trainable and pluggable virtual tokens between image features and enhanced prompt texts, bridging the gap between the encoded image features and input augmented texts. During inference, the fine-tuned virtual token embeddings are added to the original vocabulary, making it a plug-and-play method.
@@ 5.3.1. Auxiliary supervision
Visual Supervision.
Accordingly, the work of (Chen et al., 2023b) constructs a fine-grained vision instruction dataset based on Panoptic Scene Graph (PSG), called Relation-Associated Instruction (RAI-30k). In addition to standard dialogues, each instruction in RAI-30k is associated with a relation annotation in PSG, which includes mask annotations for related instances. With these additional annotations, it further supervises MLLMs with mask prediction loss using a state-of-the-art expert vision model, SAM (Kirillov et al., 2023), guiding MLLMs to focus on highly-related image content. With the additional supervision from the mask prediction loss, MLLMs are encouraged to extract features that can better represent these crucial instances, thus generating more accurate responses and mitigating vision hallucination. The intuitive idea of supervising MLLMs with grounding shows promising performance in mitigating hallucination.
Contrastive Learning.
Another line of work resort to contrastive learning to encourage better alignment at the embedding space (Jiang et al., 2023) or surpass hallucinatory response (Sarkar et al., 2025) during training.
As introduced earlier, popular MLLMs typically project the encoded vision features into the input space of a specific LLM. HACL (Jiang et al., 2023) argues that an ideal projection should blend the distribution of visual and textual embeddings. However, despite visual projection, a significant modality gap exists between textual and visual tokens, suggesting that the current learned interfaces are not effective in mapping visual representations into the textual representation space of LLMs. This issue potentially exacerbates the tendency for MLLMs to generate more hallucinations. Therefore, HACL proposes enhancing the alignment between visual and textual representations through contrastive loss. Texts with hallucinations are used as hard negative examples for image anchors. The loss pulls representations of non-hallucinating text and visual samples closer while pushing representations of non-hallucinating and hallucinative text apart.
HALVA (Sarkar et al., 2025) utilizes the philosophy of contrastive learning in the language generation process. Specifically, it first uses generative data augmentation to construct a training set of ‘hallucinated’ and ‘correct’ response pairs, by selectively altering the ground-truth phrases in the correct responses, while keeping the overall structure intact. Then, a phrase-level alignment loss is proposed to finetune the MLLM on the constructed correct and hallucinated response pairs. The data and loss provides a fine-grained supervision on phrase-level, effectively reducing hallucination.
VISTA (Li et al., 2025a) address the problem from the perspective of visual information loss in MLLMs. It introduces two complementary modules: Visual Steering Vector (VSV) that counteracts gradual visual information loss by extracting and reinforcing visual cues in activation space, and Self-Logits Augmentation (SLA) which utilizes early excitation patterns to prioritize semantically meaningful tokens. In implementation, it involves a positive prompt and a negative prompt (without image tokens), forming a contrastive-style learning paradigm.
Others.
@@
Recently, the work of CCA (Xing et al., 2024a) discusses a novel perspective regarding positional dependency modeling. The authors observe that due to the long-term decay in RoPE, LVLMs tend to hallucinate more when relevant visual cues are distant from instruction tokens in the multimodal input sequence. Therefore, they propose Concentric Causal Attention (CCA), a positional alignment strategy that mitigates the impact of RoPE long-term decay in LVLMs by naturally reducing relative distance between visual and instruction tokens.
@@ 5.3.2. Reinforcement Learning
POVID (Zhou et al., 2024a), challenges the assumption underlying previous DPO-based methods. These methods rely on the traditional preference data generation process in LLMs, where both preferred and dispreferred responses may potentially be incorrect. Therefore, this work proposes the Preference Optimization in VLLM with AI-Generated Dispreferences (POVID) framework, aiming to exclusively generate dispreferred feedback data using AI models. The dispreferred data is generated by: 1) utilizing GPT-4V to introduce plausible hallucinations into the answer, and 2) provoking inherent hallucination by introducing noise into MLLMs. In the DPO optimization framework, the ground-truth multimodal instructions serves as the preferred answers.
RLAIF-V (Yu et al., 2024) argues that the most existing RLAIF frameworks rely on expensive proprietary models, limiting the scalability. To bridge the gap, this work proposes a solution to utilize fully open-sourced MLLM to generate high-quality feedback. CLIP-DPO (Ouali et al., 2024) takes one step further by getting rid of offline data collection. Instead, starting from the initial pool of supervised fine-tuning data, it generates a diverse set of predictions, which are ranked based on their CLIP image-text similarities, and then filtered using a robust rule-based approach to obtain a set of positive and negative pairs for DPO-based training.
Following the path of RLAIF, subsequent works, including FGAIF (Jing and Du, 2024) and HSA-DPO (Xiao et al., 2024) further enhance the model by proposing fine-grained reward (feedback), including multiple types (object, attribute, relation) and segment-level reward, etc. V-DPO (Xie et al., 2024a) devises a vision-guided direct preference optimization with a synthetic dataset containing both response-contrast and image-contrast preference pairs. HDPO (Fu et al., 2024) develop three types of preference pair data targeting the following causes of MLLM hallucinations: 1) insufficient visual capabilities; 2) long context generation, and 3) multimodal conflicts. OPA-DPO (Yang et al., 2025) identifies a crucial factor of DPO: outcomes are largely contingent on whether the constructed data aligns on-policy w.r.t the initial (reference) policy of DPO. To alleviate the problems, it proposes OnPolicy Alignment (OPA)-DPO framework, which uniquely leverages expert feedback to correct hallucinated responses and aligns both the original and expert-revised responses in an on-policy manner.
@@
HalDetect (Gunjal et al., 2024) first introduces the M-HalDetect dataset for detecting hallucinations, which covers a wide range of hallucinatory content, including non-existent objects, unfaithful descriptions, and inaccurate relationships. It then proposes a multimodal reward model to detect hallucinations generated by MLLMs. The reward model is trained on the M-HalDetect dataset to identify hallucinations in the generated text. To utilize the trained reward model to reduce hallucinations, the authors introduced Fine-grained Direct Preference Optimization (FDPO). FDPO uses fine-grained preferences from individual examples to directly reduce hallucinations in generated text by enhancing the model’s ability to distinguish between accurate and inaccurate descriptions.
@@
Another similar work, ViGoR (Yan et al., 2024a), also designs a fine-grained reward model to update pre-trained MLLMs, aiming to improve visual grounding and reduce hallucination. The reward modeling in this work encompasses both human preferences and automatic metrics. Specifically, it collects human judgment and preferences for the responses generated by MLLMs by asking crowd-workers to provide fine-grained feedback at the sentence level. The collected human preference data is used to train a reward model. Additionally, it leverages advanced vision perception models to automatically score the grounding and fidelity of the text generated by an MLLM. Both sources are combined into a single reward score during the reinforcement learning procedure.
Visual Generative Feedback.
The advancement of text-to-image (T2I) generation models (Rombach et al., 2022) offers new opportunities in mitigating hallucination. ESREAL (Kim et al., 2024c) proposes to utilize a T2I model to semantically reconstruct an image from the generated caption. After that, the semantic misalignment between the two images can serve as feedback to optimize the model. In the work of ESREAL (Kim et al., 2024c), the optimization is achieved through proximal policy optimization (Schulman et al., 2017). Such visual generative feedback has also been utilized during inference time to provide visual contrast signal for decoding, as explored in ConVis (Park et al., 2024).
Another relevant work (Zhang et al., 2025b) introduces self-correcting Decoding with Generative Feedback (DeGF), a training-free algorithm that incorporates feedback from text-to-image generative models into the decoding process to effectively mitigate hallucinations in MLLMs. Specifically, DeGF generates an image from the initial response produced by MLLMs, which acts as an auxiliary visual reference and provides self-feedback to verify and correct the initial response through complementary or contrastive decoding. This is built based on the assumption: given a visual input and a textual prompt to an MLLM, if the generated response conditioned on the original image is accurate and non-hallucinatory, a text-to-image generative model should be capable of reversing this process to produce a similar image from that response.
@@ 5.3.3. Unlearning
Unlearning refers to a technique designed to induce a model to ’forget’ specific behaviors or data, primarily through the application of gradient ascent methods (Cao and Yang, 2015). Recently, unlearning for LLMs has been receiving increasing attention (Jang et al., 2022), effectively eliminating privacy vulnerabilities in LLMs. In the context of MLLMs, a recent work (Xing et al., 2024b) introduces the Efficient Fine-grained Unlearning Framework (EFUF), applying an unlearning framework to address the hallucination problem. Specifically, it utilizes the CLIP model to construct a dataset comprised of both positive samples and negative (hallucinated) samples. The training loss is applied separately for positive and negative at the sub-sentence level. To the best of our knowledge, EFUF (Xing et al., 2024b) is the first and only work that applies the unlearning framework to the task of hallucination mitigation, opening up a new path for future research.
@@ 5.4.1. Generation Intervention
Contrastive decoding (Li et al., 2022; Chuang et al., 2023) has been a popular technique in large language models to reduce hallucination and promote the plausibility of the generated text. In the language domain, the differences between a large ”expert” language model (LM) and a smaller ”amateur” LM are leveraged. The core idea is to identify and prefer outputs where the expert model’s confidence significantly exceeds that of the amateur model.
@@
In the multimodal domain, instead of using ”amateur” models, there are efforts on building a normal distribution and a ”distorted” distribution, and then conduct contrastive decoding. One representative work is VCD (Visual Contrastive Decoding) (Leng et al., 2023), which is designed to suppress the statistical biases and language priors in MLLMs during the decoding phase. The main assumption of VCD is that a distorted visual input would lead to text responses with more biases and priors. Thus, by contrasting output distributions derived from original and distorted visual inputs, VCD aims to effectively reduce the over-reliance on statistical bias and language priors. Specifically, the decoding probability distribution is calibrated using the reference (distorted) distribution.
There are many subsequent works on improving contrastive decoding to reduce hallucination. Thus, we do not dive into the details of each specific work. In the work of ICD (Wang et al., 2024f), the distorted distribution is achieved by manipulating the text instruction prompt. Specifically, role prefixes are appended to the instructions to form disturbance instructions, which can exacerbate hallucinations, as observed by the authors. With the two distributions, contrastive decoding can be performed to reduce hallucination. HIO (Chen et al., 2024a) proposes a Hallucination-Induced Optimization strategy that seeks to amplify the contrast between hallucinatory and output tokens relying on an ‘Evil-MLLM’ intentionally producing hallucination. CODE (Kim et al., 2024d) utilizes the comprehensive descriptions from model itself as visual counterpart to build contrastive decoding, aiming to correct and improve response alignment with actual visual content. VACoDe (Kim et al., 2024a) explores how to utilize multiple image augmentations in contrastive decoding and proposes to adaptively selects the augmentation with the highest contrast for each task using the proposed softmax distance metric. SID (Huo et al., 2024) aims to induce hallucination by proposing token-level disturbances. This strategy preserves only the least important vision tokens after the early decoder layers, thereby adaptively amplify vision-and-text association hallucinations during auto-regressive decoding. This strategy ensures that multimodal knowledge absorbed in the early decoder layers induces multimodal contextual rather than aimless hallucinations. DAMRO (Gong et al., 2024) takes inspiration from the analysis of the attention pattern in the vision encoder and MLLM, and finds that the distributions tend to focus on particular background tokens rather than the referred objects in the image, potentially introducing hallucination. To address the issue, it proposes to employ classification token (CLS) of ViT to filter out high-attention outlier tokens scattered in the background and then eliminate their influence by contrastive decoding. VaLiD (Wang et al., 2024c) utilizes the technique of contrastive decoding from a vision encoder perspective. It first identifies that hallucinations in MLLMs may originate at the initial stage of visual perception in vision encoder. Based on this, it utilizes information from the distorted visual layers and applies contrastive decoding. CMVED (Li et al., 2025b), when build contrastive distribution, performs distortion by selectively masking the value vectors associated with high cross-modal attention weights in self-attention layers, which suppresses important inter-modality correlations while retaining the spurious ones in distorted distribution. VASparse (Zhuang et al., 2025) levearges the observation of sparse activation of attention in MLLMs and designs a token sparsification strategy that balances efficiency and trustworthiness. Based on that, it further introduces a sparse-based visual contrastive decoding method to recalibrate the distribution of hallucinated outputs. Octopus (Suo et al., 2025) argues that the mechanism of hallucination occurrence is a complex hybrid and different samples (or tokens), while current contrastive decoding methods mostly apply the same disturbed manner for all samples and generative steps. Thus, it focuses on guiding the model to dynamically organize the contrastive decoding workflow and selecting the appropriate contrastive decoding strategy based on different inputs.
@@
MARINE (Zhao et al., 2024a) proposes a training-free approach. It employs an additional vision encoder for object grounding and utilizes the grounded objects to guide the decoding process. Specifically, it innovatively adapts the classifier-free guidance (Ho and Salimans, 2022) technique to implement guided decoding, showing promising performance in emphasizing the detected objects while reducing hallucination in the text response.
@@
HALC (Chen et al., 2024e) provides a key insight that when decoding a specific token in the MLLM, identifying a token-wise optimal visual context to provide the most informative visual grounding can effectively reduce hallucination. Visual context refers to the visual tokens that can be grounded from the generated text response. An oracle study showed that decoding from the provided optimal visual contexts eliminates over 84.5% of hallucinations. Based on the insight and observation, the authors designed mechanisms to locate the fine-grained visual information to correct each generated token that might be hallucinating. This is essentially a visual content-guided decoding strategy. In addition to token-level correction, HALC also incorporates a matching-based beam search that utilizes a visual matching score to steer the generation of the final outputs, balancing both object hallucination mitigation and text generation quality.
The work of DeCo (Wang et al., 2024b) has conducted an empirical analysis and find that MLLMs are not blind; they can recognize objects in the preceding layers, but this recognition is suppressed in later layers, leading to hallucinations. Based on this observation, it proposes Dynamic Correction Decoding with preCeding-Layer Knowledge (DeCo) to mitigate hallucinations for MLLMs. It dynamically selects preceding layer and utilizes its prior knowledge as a guidance to correct the final output logits.
SumGD (Min et al., 2024) aims to address hallucination caused by language prior. The approach employs a summarization technique that selectively retains essential information from previously generated sentences, encouraging LVLMs to more effectively incorporate image information. To minimize unnecessary intervention for preserving text quality, the summarization is referenced only when predicting image-related POS tokens, which require image-specific details.
Visual Amplification
Many works (Favero et al., 2024) have identified that one cause of hallucination is the weak dependency on the visual prompt during generation. And this dependency even decreases as more tokens are generated (Favero et al., 2024), which is consistent with the observation of longer output tend to generate more hallucination in the work of (Han et al., 2024a). To address this, M3ID (Favero et al., 2024) proposes a novel decoding algorithm, which models two output distributions with and without the image prompt as condition. During generation, the final output is determined together by two distributions, explicitly amplifying the image prompt.
IBD (Zhu et al., 2024) proposes an image-biased decoding strategy, which is also closely related to contrastive decoding. Specifically, IBD involves computing a more reliable next-token probability distribution by contrasting the predictions of the original model with those of an image-biased model, which focuses more on the image information. The image-based model is created by modifying the attention weight matrix structure within the original model, without altering its parameters. This approach emphasizes the knowledge of the image-biased model and diminishes that of the original model, which may be text-biased. Thus, it encourages the extraction of correct content while suppressing hallucinations resulting from textual over-reliance.
AGLA (An et al., 2024) argues that current MLLMs demonstrate attention deficiency that causes the ignorance of prompt-relevant object features in images. Therefore, it introduces prompt-relevant local attention to generate an augmented view of the image for hallucination mitigation. Specifically, it design an image-prompt matching scheme that computes the prompt-relevant attention and further generates an augmented view of the original image, better supporting object-related queries.
RVD (Zhong et al., 2024)(Residual Visual Decoding) proposes a straightforward strategy. By residual connecting the visual information and the current user instruction, distributions that emphasizing the visual information are derived to revise the original output distribution. This strategy has shown effectiveness on amplifying the visual information.
Another representative work PAI (Liu et al., 2024d) (Pay Attention to Image) intervenes in the inference process to make it more image-centric, following the original image perception direction. To achieve this, the method focuses on the self-attention heads in the decoder layers of MLLMs. It enhances the attention weights for image tokens in their original directions during inference. This allows the model to use the updated attention matrix to calculate the hidden states for the generated token, thereby incorporating more consideration for image representation during the generation process.
EAH (Zhang et al., 2024b) also identifies this problem of visual attention sink. To address this, it proposes a training-free method named Enhancing Attention Heads (EAH), an approach designed to enhance the convergence of image tokens attention sinks in the shallow layers. EAH identifies the attention head that shows the vision sink in a shallow layer and extracts its attention matrix. This attention map is then broadcast to other heads in the layer, thereby strengthening the layer to pay more attention to the image itself. A similar idea has also been explored in VHR (He et al., 2024b), a training-free approach aimed at enhancing the model’s reliance on visual context rather than language priors. This achieved by first identifying key attention heads based on their VHD scores and then amplifying their contributions during generation.
PM (Mao et al., 2025) proposes an intuitive strategy that adaptively modulates spatial resolution, processing key visual regions at higher detail while compressing less relevant areas. It refines the visual input by magnifying relevant regions while compressing less critical areas, implemented as a resampling process.
@@
The work of OPEAR (Huang et al., 2023b) makes an interesting observation that most hallucinations are closely tied to the knowledge aggregation patterns manifested in the self-attention matrix, i.e., MLLMs tend to generate new tokens by focusing on a few summary tokens rather than all the previous tokens. Such a partial over-trust inclination results in neglecting image tokens and describing the image content with hallucination. Based on this observation, a decoding method for MLLMs grounded in an Over-trust Penalty and a Retrospection-Allocation strategy is proposed. First, a penalty term on the model logits is introduced during the MLLM beam-search decoding process to mitigate the over-trust issue. Additionally, to handle the hard cases that cannot be addressed by the penalty term, a more aggressive strategy called the rollback strategy is proposed to retrospect the presence of summary tokens in the previously generated tokens and reallocate the token selection if necessary. A follow-up work HELPD (Yuan et al., 2024) modified the penalty by incorporating visual attention into the penalty score computation and makes the final logits place more emphasis on the image input, amplifying the visual attention as discussed above.